Objectives only

  1. Suppose you open the Economic Times and see that 30 year Treasury bonds are yielding 1.6%. This is the
  1. risk premium
  2. real return
  3. nominal return
  4. inflation premium
  5. apocalypse

2. Coke issued 7% coupon bonds at par value one year ago. Immediately after issuance of these bonds, market interest rates decreased to 6.5%. Which of the statements below is true regarding these Coke bonds?

  1. The bonds became premium bonds when market rates decreased to 6.5%.
  2. The bonds became discount bonds when market rates decreased to 6.5%.
  3. The coupon interest payment on the bonds adjusted to the market rates when market interest rates decreased.
  4. Answers a and c are true.
  5. Answers b and c are true.

3. Which of the statements below is true about the cash flows from stocks and bonds?

  1. In order to remain solvent, public corporations must pay interest coupons, principal amounts that are due, and common stock dividends every period.
  2. Dividends and coupon interest payments are optional for public corporations.
  3. Dividends are optional but interest coupons and principal amount due must be paid, if a firm is to remain solvent.
  4. Common stock dividends may be delayed but must be paid eventually, while coupon interest payments must be paid by their due date.

4. Suppose you are using the PI method to evaluate independent capital budgeting projects. You should choose the project with the highest PI, as long as that PI is greater than zero.

  1. true
  2. false

5. ABC Co. bonds sell for $950. The coupon interest rate on these 8 year bonds is 6%. However, interest is paid semiannually. The par value is $1,000. If they are purchased at the market price, what is the yield to maturity?

a. 3.41%

b. 6.82%

c. 2.75%

d. 5.49%

e. 6.12%

6.Dairy Co. bonds have a 9% annual coupon rate, a face value of $1,000 and mature in 7 years. Interest is paid semi-annually. The current market yield on bonds in this risk class is 8.5%. Calculate the current market price of this bond.

a. $ 984

b. $1,040

c. $ 986

d. $1,059

e. $1,026
Answer #1

1.
Nominal return

2.
If rates decrease to less than coupon rate, bonds sell for more than par value i.e., sell for premium
The bonds became premium bonds when market rates decreased to 6.5%.

3.
Dividends are optional but interest coupons and principal amount due must be paid, if a firm is to remain solvent.

The company ahs a choice regarding dividends but copulsoriyl pay bondholders

4.
False, PI should be greater than 1

5.
Using financial calculator
N=8*2
PV=-950
PMT=6%*1000/2
FV=1000
CPT I/Y=3.41%

Hence, yield to maturity=3.41%*2=6.82%

6.
Using financial calculator
N=7*2
I/Y=8.5%/2
PMT=-9%*1000/2
FV=-1000
CPT PV=1025.98